OpenStudy (anonymous):

What is the equation in standard form of the line which passes through (1,-3) and has a slope of 2 2x-y=5 2x-y=-5 2x+y=5 2x+y=-5

OpenStudy (ivettef365):

use formula y = mx + b, you have m=2 now substitute x = 1 and y = -3 you have -3 = 2(1) + b find the value of b

OpenStudy (anonymous):

So it's the last one

OpenStudy (ivettef365):

well you have b = -5 then you have the equation y = 2x - 5 let see which one is the same, you can switch numbers around as long as you change the sign

OpenStudy (ivettef365):

y = 2x -5 if I switch the y and the -5 then I have 5 = 2x - y

OpenStudy (ivettef365):

which is the same as 2x - y = 5

OpenStudy (anonymous):

Can you help me with another problem please. What is the equation of the line in standard form that passes through the point (-1,4) and is parallel to the line y=2x-3

OpenStudy (anonymous):

X-2y=6 x+2y=6 2x-y=-6 2x+y=-6

OpenStudy (ivettef365):

If it is parallel it means that the slope is the same you have the equation y = 2x -3, the slope is 2, since formula is y = mx + b where m = slope do you see this?

OpenStudy (anonymous):

So it's the last one

OpenStudy (ivettef365):

then you have the points (-1,4), to find the equation for this point substitute the x = -1 and y = 4, you need to find value of b then you have 4 = 2(-1) + b what is the value of b ??

OpenStudy (anonymous):

6 just tell me the answer

OpenStudy (ivettef365):

ok you have y = 2x + 6 when you switch the y and the 6 you have -6 = 2x -y which is the same as 2x - y = -6